# Masked Autoencoder Self Pre-Training for Defect Detection in Microelectronics

### v1

@@ Abstract
Whereas in general computer vision, transformer-based architectures have quickly become the gold standard, microelectronics defect detection still heavily relies on convolutional neural networks (CNNs). We hypothesize that this is due to the fact that a) transformers have an increased need for data and b) labelled image generation procedures for microelectronics are costly, and labelled data is therefore sparse. Whereas in other domains, pre-training on large natural image datasets can mitigate this problem, in microelectronics transfer learning is hindered due to the dissimilarity of domain data and natural images. Therefore, we evaluate self pre-training, where models are pre-trained on the target dataset, rather than another dataset. We propose a vision transformer (ViT) pre-training framework for defect detection in microelectronics based on masked autoencoders (MAE). In MAE, a large share of image patches is masked and reconstructed by the model during pre-training. We perform pre-training and defect detection using a dataset of less than 10.000 scanning acoustic microscopy (SAM) images labelled using transient thermal analysis (TTA). Our experimental results show that our approach leads to substantial performance gains compared to a) supervised ViT, b) ViT pre-trained on natural image datasets, and c) state-of-the-art CNN-based defect detection models used in the literature. Additionally, interpretability analysis reveals that our self pre-trained models, in comparison to ViT baselines, correctly focus on defect-relevant features such as cracks in the solder material. This demonstrates that our approach yields fault-specific feature representations, making our self pre-trained models viable for real-world defect detection in microelectronics.
@@
Masked Autoencoder Vision Transformer Pretraining Microelectronics LED Solder Joint Quality Control
@@ 1 Introduction
Figure 1: Masked Autoencoder Framework. In MAE pre-training, a large share (e.g. $75\%$) of input patches is masked. The encoder (ENC) then performs self-attention for visible patches only. After adding mask-tokens, which serve as placeholders for missing inputs, the decoder (DEC) reconstructs the complete image. Thereby, the model learns meaningful representations of input data.
@@
In particular, deep learning based vision models have recently shown strong performance in AOI tasks like defect detection [41, 42, 27, 18]. While convolutional neural networks (CNNs) have traditionally been the dominant choice for deep learning based computer vision, vision transformer (ViT) models have recently emerged as the gold standard in the field [20, 22, 35]. However, the application of transformer models to defect detection in microelectronics remains underexplored, with the field often relying on conventional CNNs and recent survey studies sometimes not mentioning transformer models at all [27, 18]. One main reason for this could be that training transformers from scratch usually requires a large amount of data [40, 13, 46]. In microelectronics manufacturing - especially for microscale solder joints - labelled image data collection often relies on expensive procedures, resulting in small and imbalanced datasets [1]. Additionally, fine-tuning transformers pre-trained on large datasets usually requires at least some similarity between the pre-training domain and the target domain [2]. This circumstance is problematic for microelectronics since most pre-trained transformers are trained on natural image datasets like ImageNet [11] that are highly dissimilar from target datasets in this domain (see Figure 2).
@@
Inspired by other domains facing similar challenges, such as healthcare, this paper explores the potential of a pure transformer model for microelectronics defect detection by leveraging self pre-training, where models are pre-trained directly on the target dataset rather than an extensive natural image dataset [50]. Specifically, we use masked autoencoders (MAEs) [16], which mask a large share of image patches and task the model with reconstructing the missing inputs (see Figure 1). MAEs are resource-efficient and well-suited for smaller datasets, as they do not require large batch sizes like contrastive pre-training approaches. Additionally, the repeated randomized masking of different image patches presents the autoencoder with diverse reconstruction tasks, even with limited data. Using a dataset of less than 10.000 scanning acoustic microscopy (SAM) images of microscale LED solder joints labelled using transient thermal analysis (TTA), we compare our approach to a) purely supervised ViT, b) ViT pre-trained on a natural image dataset, and c) state-of-the-art CNN architectures proposed in the literature on industrial defect detection. In particular, we fine-tune our models for a regression task where models predict how far a given LED is from failure.
1.0.1 Contributions
Our experimental findings indicate that, with our approach, models learn generalizable representations in self pre-training that yield substantial performance increases in defect detection for microelectronics. In particular, our model outperforms a) purely supervised ViTs by $21.1\%$, b) ViTs pre-trained on Imagenet by $10.2\%$ and c) state-of-the-art CNN defect detection models by $5.3\%$. Also, we find that our models learn fault-specific representations, focusing on actual damages in the solder material when performing defect detection, while other ViT models have scattered attention or learn shortcuts. In addition, due to the simplicity and low data requirements of our framework, it can be easily integrated into real-world inline quality control procedures such as AOI.
@@ 2.0.1 Vision Transformer
ViT [13] is a vision adaption of the self-attention based transformer architecture [40], which constitutes the backbone of large language models. ViT has demonstrated effectiveness across various vision tasks, including image classification, object detection, and segmentation [19, 33, 37]. Related research on ViT includes domain-specific fine-tuning of foundational models and architectural adaptions that aim at improved domain generalization or increased data efficiency: Alijiani et al. [2] discuss various ViT domain adaption and generalization approaches on the feature extraction level, the instance level and the model architecture level. Papa et al. [26] compare ViT approaches aiming at increased data efficiency, such as Data Efficient Transformers (DeiT) [38, 39] or Shifted Window Transformers (Swin) [21]. Since we leverage representations learned in pre-training, our approach is similar to domain specific fine-tuning. However, our framework differs from such approaches since we pre-train ViT from scratch on the target dataset, rather than using foundational models pre-trained on large natural image datasets.
@@ 2.0.2 Self-Supervised Learning
SSL aims to learn meaningful representations from data without labels. Instead of labels, intrinsic relationships of the data serve as guidance for the training process. Balestriero et al. [3] distinguish four sub-types of SSL approaches: (1) deep metric learning approaches such as SimCLR [10] or NNCLR [14], where models recognize similarity in differently transformed inputs, (2) self-distillation learning approaches like BYOL [15], SimSIAM [9] or DINO [7], where different input transforms serve as inputs into two separate encoders before letting one serve as a predictor of the other’s output, (3) canonical correlation analysis approaches such as VICReg [5], BarlowTwins [45] or SWAV [6], which are founded on the principle of understanding relationships between two variables by examining their cross-covariance matrices and (4) masked image modelling (MIM) approaches like SimMIM [43], where images undergo partial masking and models reconstruct the missing inputs. We follow the MIM approach due to its simplicity, efficacy, and low data requirements.
[…]
@@ 2.0.3 Microelectronics Quality Control
In microelectronics research and manufacturing, image-based quality control procedures like AOI are standard practice [1]. Deep learning based approaches have significantly advanced the field in recent years, including but not limited to the subdomain of solder joints [41, 42, 27, 18]. Samavatian et al. [29] present an iterative machine learning framework that enhances the accuracy of solder joint lifetime prediction by utilizing a self-healing dataset iteratively injected into a correlation-driven neural network (CDNN). Salameh et al. [28] demonstrate the use of deep neural networks, combined with finite element simulations, as a rapid and comprehensive tool for solder joint reliability analysis under mechanical loading. Muench et al. [24] propose a methodology for predicting damage progression in solder contacts and compare an MLP network with a long short-term memory (LSTM) model using production-like synthetic data. Zhang et al. [48] employ various deep learning models, including CNN and LSTM models, for automatic solder joint defect detection using X-ray images. Zhang et al. [49] combine CNN and transformer components in a model intended for printed control board solder joints. However, Zhang et al. use transformer blocks only as small parts of their network aided by convolutional layers and report poor performances for pure transformer models [49]. We believe that this is due to a lack of domain-specific pre-training. Our approach, in contrast, demonstrates the potential of pure transformer architectures for defect detection without architectural crutches like CNN layers.
@@ 3 Dataset
We use scanning acoustic microscopy (SAM) images and transient thermal analysis (TTA) measurements from a dataset of microscale solder joints of high-power LEDs [31], which is publicly available on Kaggle [30]. The dataset contains SAM and TTA data for 1.800 LEDs at 5 points of time, with nine LED types and five lead-free solder pastes. To simulate ageing, LED panels underwent thermal shock cycles (TSC) from $-40$°C to $125$°C, a standard quality control process to evaluate the reliability and durability of solder joints under extreme conditions. We use a train, validation and test split of 60%, 20% and 20% respectively. All reported metrics and all shown example outputs refer to the test dataset.
Table 1: Dataset. Repeated cooling and heating of LED solder joints simulates aging through thermomechanical fatigue. Over time, cracks gradually emerge, impeding the heat flow through the solder material. $\Delta B_{max}$ values indicate relative degradation and LEDs with $\Delta B_{max}>20\%$ are classified as defective ()(*)( ∗ ).
TSC 0 100 500 1.000 1.500
@@ 3.1 Scanning Acoustic Microscopy (SAM)
We use SAM images for pre-training and as model inputs in downstream tasks. SAM imaging offers detailed information on material interfaces, allowing the detection of critical defects such as voids, cracks, and delaminations within solder joints [23]. SAM images were taken after 0, 100, 500, 1.000, and 1.500 TSC, resulting in 5 images per LED and 9.000 images in total (see Table 1). The original SAM images contain 2 LEDs and are cropped so that each image shows an individual solder joint. All images are resized to a $64\times 64\times 1$ pixel format. For each LED image, bright boundary regions depict gaps between the darker solder pads. On the solder pads, small circles indicate voids in the solder material, from which cracks can emerge over time. Bright, non-circular structures inside the pads are cracks in the solder material (see Table 1).
@@ 4.1.1 Vision Transformer
patches. After flattening and applying a linear mapping $E\in\mathbb{R}^{P^{2}\times D}$ to project the patch into the embedding space $\mathbb{R}^{D}$, positional embeddings $E_{pos}\in\mathbb{R}^{N\times D}$ are added to the elements of the sequence to retain the spatial location of each patch embedding. Lastly, a learnable class token $x_{class}$ is prepended to the sequence to store global information during processing.Images are thus transformed into a sequence $X$ of $N+1$ embeddings, which serve as tokens for MSA:
@@ 4.1.2 Masked Autoencoder Pre-Training
For MAE pre-training, a ViT encoder combined with a lightweight transformer decoder is tasked with reconstructing missing image patches (see Figure 1). Depending on the masking share $S$, which is usually $0.75$, a random subset $M$ of $\lfloor N\cdot S\rfloor$ indexes is sampled from the set of indexes $I=\{1,\dots,N\}$ without replacement. $X$ is split into a set of masked patches $X_{M}$ and a set of visible patches $X_{I\setminus M}$. Note that $X_{M}$ and $X_{I\setminus M}$ are absolute complement with respect to $X_{I}$, i.e. $X_{m}\cup X_{I\setminus M}=X_{I}$ and $X_{m}\cap X_{I\setminus M}=\emptyset$. After masking, the sequence consists of $\lfloor N\cdot S\rfloor$ embeddings:
@@
Inputs are then fed into the autoencoder network. The MAE encoder operates only on the visible patches, i.e. on a $1-S$ share of the total input patches. The decoder, in contrast, receives as input a full sequence of tokens including a) the embeddings of visible patches and b) mask tokens, which are shared and learnable embeddings serving as placeholders for masked patches, similar to BERT [12]. The model minimizes the pixel-wise MSE loss for masked patches with respect to its parameters $\theta$. The loss is computed for masked patches only, i.e. for $k\in M$:
@@ 5.1 Hyperparameter Tuning
Because of the different structure of SAM images and due to the smaller image size compared to standard datasets like ImageNet, we hypothesize that smaller patch sizes, different augmentations, and different mask ratios than those used for standard ViT could achieve optimal performance. To evaluate hyperparameter settings, we pre-train models with MAE for 200 epochs on our dataset, and fine-tune for defect detection for 100 epochs.
@@
For augmentations, a combination of horizontal and vertical flipping and fixed-size cropping achieved the best results, while for ImageNet horizontal flipping and random-sized cropping are used [13]. In contrast to natural images, vertical flipping is a valid augmentation for microelectronics, where image orientation is not relevant. Also, we suspect that fixed-size cropping works best since random-sized crops of $64\times 64$ images suffer greater quality loss by resizing operations than the $224\times 224$ images in standard large datasets like ImageNet.
@@ 5.2 MAE Self Pre-Training
We conduct MAE pre-training on the target SAM image dataset, utilizing the hyperparameters optimized through our previous studies. MAE with ViT encoders of sizes Ti, S, and B are pre-trained for 1.000 epochs on two L40S GPUs respectively. In Figure 3, we plot the MSE reconstruction loss for masked input patches $\mathcal{L}_{mae}$ for each model size. Notably, the S-sized encoder achieved the worst performance in reconstruction, although having more parameters than the Ti-sized encoder. However, the even larger B-sized encoder significantly outperformed all smaller sizes. We take this to indicate that a) for smaller images, even very small encoders like ViT-Ti can achieve good results and that b) there is a non-linear relationship between model capacity and performance. For smaller datasets like ours, there appears to be a critical threshold where increased model complexity transitions from potentially harmful (due to overfitting) to beneficial. Although ViT-B achieved the best results overall, ViT-Ti could be the model of choice in resource-critical applications.
@@
(b) 1500 TSC
Figure 4: MAE Reconstructions. For (a) TSC 0 and (b) TSC 1500, original images, masked inputs, and the model’s reconstructions are depicted.

### v2

@@ Abstract
While transformers have surpassed convolutional neural networks (CNNs) in various computer vision tasks, microelectronics defect detection still largely relies on CNNs. We hypothesize that this gap is due to the fact that a) transformers have an increased need for data and b) (labelled) image generation procedures for microelectronics are costly, and data is therefore sparse. Whereas in other domains, pre-training on large natural image datasets can mitigate this problem, in microelectronics transfer learning is hindered due to the dissimilarity of domain data and natural images. We address this challenge through self pre-training, where models are pre-trained directly on the target dataset, rather than another dataset. We propose a resource-efficient vision transformer (ViT) pre-training framework for defect detection in microelectronics based on masked autoencoders (MAE). We perform pre-training and defect detection using a dataset of less than 10,000 scanning acoustic microscopy (SAM) images. Our experimental results show that our approach leads to substantial performance gains compared to a) supervised ViT, b) ViT pre-trained on natural image datasets, and c) state-of-the-art CNN-based defect detection models used in microelectronics. Additionally, interpretability analysis reveals that our self pre-trained models attend to defect-relevant features such as cracks in the solder material, while baseline models often attend to spurious patterns. This shows that our approach yields defect-specific feature representations, resulting in more interpretable and generalizable transformer models for this data-sparse domain.
@@
Masked Autoencoder Vision Transformer Pretraining Self-Supervised Learning Data-Efficient Learning Microelectronics
@@ 1 Introduction
Figure 1: Masked Autoencoder for Microelectronics. In MAE pre-training, a large share of input patches is masked. The encoder (ENC) then performs self-attention for visible patches only. After adding mask-tokens, which serve as placeholders for missing inputs, the decoder (DEC) reconstructs the complete image. Thereby, the model learns meaningful representations of input data.
@@
Recently, vision transformer (ViT) models have emerged as the gold standard in computer vision [20, 21]. However, the application of transformer models to defect detection in microelectronics remains underexplored, with the field largely relying on convolutional neural networks (CNN) and recent survey studies sometimes not mentioning transformer models at all [25, 19]. One main reason for this could be that training transformers from scratch usually requires a large amount of data [32, 14, 36]. In microelectronics manufacturing - especially for microscale solder joints - labelled image data collection often relies on expensive procedures, resulting in small and imbalanced datasets [1]. Additionally, fine-tuning transformers pre-trained on large datasets usually requires at least some similarity between the pre-training domain and the target domain [2]. This poses a challenge for microelectronics since most pre-trained transformers are trained on natural image datasets like ImageNet [12] that are highly dissimilar from target datasets in this domain (see Figure 2).
@@
Inspired by other domains facing similar challenges, such as healthcare, this paper explores the potential of a pure transformer model for microelectronics defect detection by leveraging self pre-training, where models are pre-trained directly on the target dataset rather than an extensive natural image dataset [40]. Specifically, our framework is based on masked autoencoders (MAEs) [17], which mask a large share of image patches and task the model with reconstructing the missing inputs (see Figure 1). MAEs are resource-efficient and well-suited for smaller datasets, as they do not require large batch sizes like contrastive pre-training approaches. Additionally, the repeated randomized masking of different image patches presents the autoencoder with diverse reconstruction tasks, even with limited data. Using a dataset of less than 10,000 scanning acoustic microscopy (SAM) images of microscale LED solder joints labelled using transient thermal analysis (TTA), we compare our approach to a) purely supervised ViT, b) ViT pre-trained on a natural image dataset, and c) state-of-the-art CNN architectures proposed in the literature on industrial defect detection. In particular, we train our models for a regression task where models predict how far a given LED is from failure. We make the following contributions:
  • We adapt masked autoencoder pre-training to defect detection under severe data scarcity in the microelectronics domain by self pre-training models directly on the target dataset, rather than on large natural image datasets.
  • We demonstrate that this domain-specific self pre-training on less than 10,000 SAM images significantly outperforms supervised ViTs, ImageNet-pretrained ViTs, and several CNN architectures for defect detection. Our largest model improves mean squared error by up to 25%, while requiring less than 12 hours of GPU time on a single A100 for both pre-training and fine-tuning, enabling resource-efficient application of vision transformers in this data-scarce domain.
  • We show that, compared to baseline models, our approach yields defect-specific feature representations in pure transformer architectures. Our self pre-trained models focus on meaningful defect regions, leading to improved interpretability and generalizability, whereas baseline approaches often attend to irrelevant features or learn shortcuts.
@@ 2.0.1 Vision Transformer
ViT [14] is a vision adaption of the self-attention based transformer architecture [32]. Related research on ViT includes domain-specific fine-tuning of foundational models and architectural adaptions that aim at improved domain generalization or increased data efficiency: Alijiani et al. [2] discuss various ViT domain adaption and generalization approaches on the feature extraction level, the instance level and the model architecture level. Since we leverage representations learned in pre-training, our approach is similar to domain specific fine-tuning. However, our framework differs from such approaches since we pre-train ViT from scratch on the target dataset, rather than using foundational models pre-trained on large natural image datasets.
@@ 2.0.2 Self-Supervised Learning
SSL aims to learn meaningful representations from data without labels. Instead of labels, intrinsic relationships of the data serve as guidance for the training process. Balestriero et al. [3] distinguish four sub-types of SSL approaches: (1) deep metric learning approaches such as SimCLR [11] or NNCLR [15], where models recognize similarity in differently transformed inputs, (2) self-distillation learning approaches like BYOL [16], SimSIAM [10] or DINO [7], where different input transforms serve as inputs into two separate encoders before letting one serve as a predictor of the other’s output, (3) canonical correlation analysis approaches such as VICReg [5], BarlowTwins [35] or SWAV [6], which are founded on the principle of understanding relationships between two variables by examining their cross-covariance matrices and (4) masked image modelling (MIM) approaches like SimMIM [33], where images undergo partial masking and models reconstruct the missing inputs. We follow the MIM approach due to its its simplicity and effectiveness in low-data environments.
In MIM, iGPT [8] first demonstrated the potential of applying masked language modeling strategies to the vision domain by predicting masked pixels in a sequence-like manner. BEiT [4] introduced a discrete token prediction objective analogous to masked language modeling in NLP and showed that visual tokens enhance pre-training for image-related tasks. In MAE [17], an encoder receives approximately $25\%$ of the unmasked patches as input, and a lightweight decoder reconstructs the full image (see Figure 1). MAE pre-training has been demonstrated to achieve state-of-the-art performance with minimal data and compute requirements [37]. Our approach builds on the work of He et al. [17], adapting their framework to the requirements of our use case.
@@ 2.0.3 Microelectronics Quality Control
Deep learning based approaches have significantly advanced the field in recent years, including but not limited to the subdomain of solder joints [25, 19]. Samavatian et al. [27] present an iterative machine learning framework that enhances the accuracy of solder joint lifetime prediction. This is achieved by utilizing a self-healing dataset that is iteratively injected into a correlation-driven neural network (CDNN). Salameh et al. [26] demonstrate the use of deep neural networks, combined with finite element simulations, as a rapid and comprehensive tool for solder joint reliability analysis under mechanical loading. Muench et al. [23] propose a methodology for predicting damage progression in solder contacts and compare a multi-layer perceptron network with a long short-term memory (LSTM) model using production-like synthetic data. Zhang et al. [38] employ various deep learning models, including CNN and LSTM models, for automatic solder joint defect detection using X-ray images. Zhang et al. [39] combine CNN and transformer components in a model intended for printed control board solder joints. However, Zhang et al. use transformer blocks only as small parts of their network aided by convolutional layers and report poor performances for pure transformer models [39]. We believe that this is due to a lack of domain-specific pre-training. Our approach, in contrast, demonstrates the potential of pure transformer architectures for defect detection without architectural crutches like CNN layers.
@@ 3 Dataset
We use scanning acoustic microscopy (SAM) images and transient thermal analysis (TTA) measurements from a dataset of microscale solder joints of high-power LEDs [29], which is publicly available on Kaggle [28]. The dataset contains SAM and TTA data for 1,800 LEDs at 5 points of time, with nine LED types and five lead-free solder pastes. To simulate ageing, LED panels underwent thermal shock cycles (TSC) from 40-40°C to 125125°C, a standard quality control process to evaluate the reliability and durability of solder joints under extreme conditions. We use a train, validation and test split of 60%, 20% and 20% respectively. All reported metrics and all shown example outputs refer to the test dataset.
Table 1: Dataset. Repeated cooling and heating of LED solder joints simulates aging through thermomechanical fatigue. Over time, cracks gradually emerge, impeding the heat flow through the solder material. $\Delta B_{max}$ values indicate relative degradation and LEDs with $\Delta B_{max}>20\%$ are classified as defective ()(*).
TSC 0 100 500 1,000 1,500
@@ 3.1 Scanning Acoustic Microscopy (SAM)
We use SAM images for pre-training and as inputs for defect detection. SAM imaging offers detailed information on material interfaces, allowing the detection of critical defects such as voids, cracks, and delaminations within solder joints [22]. SAM images were taken after 0, 100, 500, 1,000, and 1,500 TSC, resulting in 5 images per LED and 9,000 images in total (see Table 1). The original SAM images contain 2 LEDs and are cropped so that each image shows an individual solder joint in a $64\times 64\times 1$ pixel format. For each LED image, bright boundary regions depict gaps between the darker solder pads. On the solder pads, small circles indicate voids in the solder material, from which cracks can emerge over time. Bright, non-circular structures inside the pads are cracks in the solder material (see Table 1).
@@ 4.1.1 Vision Transformer
patches. After flattening and applying a linear mapping $E\in\mathbb{R}^{P^{2}\times D}$ to project the patch into the embedding space $\mathbb{R}^{D}$, positional embeddings $E_{pos}\in\mathbb{R}^{N\times D}$ are added to the elements of the sequence to retain the spatial location of each patch embedding. Lastly, a learnable class token $x_{class}$ is prepended to the sequence to store global information relevant to the $\Delta B_{max}$ prediction during processing. Images are thus transformed into a sequence $X$ of $N+1$ embeddings, which serve as tokens for MSA:
@@ 4.1.2 Masked Autoencoder Pre-Training
For MAE pre-training, a ViT encoder combined with a lightweight transformer decoder is tasked with reconstructing missing image patches (see Figure 1). Depending on the masking share $S$, which is usually $0.75$, a random subset $M$ of $\lfloor N\cdot S\rfloor$ indexes is sampled from the set of indexes $I=\{1,\dots,N\}$ without replacement. $X$ is split into a set of masked patches $X_{M}$ and a set of visible patches $X_{I\setminus M}$. Note that $X_{M}$ is the absolute complement of $X_{I\setminus M}$ with respect to $X_{I}$, i.e. $X_{m}\cup X_{I\setminus M}=X_{I}$ and $X_{m}\cap X_{I\setminus M}=\emptyset$. After masking, the sequence consists of $\lfloor N\cdot S\rfloor$ embeddings:
@@
Inputs are then fed into the autoencoder network. The MAE encoder operates only on the visible patches, i.e. on a 1S1-S share of the total input patches. The decoder, in contrast, receives as input a full sequence of tokens including a) the embeddings of visible patches and b) mask tokens, which are shared and learnable embeddings serving as placeholders for masked patches, similar to BERT [13]. The model minimizes the pixel-wise mean squared error (MSE) loss for masked patches with respect to its parameters $\theta$. The loss is computed for masked patches only, i.e. for $k\in M$:
@@ 5.1 Hyperparameter Tuning
Because of the distinct structure of SAM images and due to the smaller image size compared to standard datasets, we hypothesize that smaller patch sizes, different augmentations, and different mask ratios than those used for standard ViT could achieve optimal performance. To evaluate hyperparameter settings, we pre-train models with MAE for 200 epochs on our dataset, followed by supervised fine-tuning for defect detection over 100 epochs. We report the MSE with respect to the target $\Delta B_{\text{max}}$ after fine-tuning.
@@
For augmentations, a combination of horizontal and vertical flipping and resized cropping using a fixed crop-size achieved the best results, while for ImageNet horizontal flipping and random-sized cropping are used [14]. In contrast to natural images, vertical flipping is a valid augmentation for microelectronics, where image orientation is not relevant. Also, we suspect that fixed-size cropping works best since random-sized crops of $64\times 64$ images suffer greater quality loss by resizing operations than the $224\times 224$ images in standard large datasets like ImageNet.
@@ 5.2 MAE Self Pre-Training
We conduct MAE pre-training on the target SAM image dataset, utilizing the hyperparameters optimized through our previous studies. MAE with ViT encoders of sizes Ti, S, and B are pre-trained for 1,000 epochs respectively. In Figure 3, we plot the MSE reconstruction loss for masked input patches $\mathcal{L}_{mae}$ for each model size. Notably, the S-sized encoder achieved the worst performance in reconstruction, although having more parameters than the Ti-sized encoder. However, the even larger B-sized encoder significantly outperformed all smaller sizes. We take this to indicate that a) for smaller images, even very small encoders like ViT-Ti can achieve good results and that b) there is a non-linear relationship between model capacity and performance. For smaller datasets like ours, there appears to be a critical threshold where increased model complexity transitions from potentially harmful (due to overfitting) to beneficial. Although ViT-B achieved the best results overall, ViT-Ti could be the model of choice in resource-critical applications.
@@
(b) 1,500 TSC
Figure 4: MAE Reconstructions. For (a) TSC 0 and (b) TSC 1,500, original images, masked inputs, and the model’s reconstructions are depicted.
Computational Effort All MAE pre-training experiments were conducted on a single A100 GPU. For our best-performing model ViT-B, the full pre-training process completes in under 8 hours, and fine-tuning for defect detection takes less than 1 hour. For the ViT-Ti model, which already outperforms all evaluated baselines on defect detection, the whole process of pre-training and fine-tuning takes under 2 hours. This highlights the computational accessibility of our approach, even in constrained environments.